OpenStudy (anonymous):

Something km =620,000 m

OpenStudy (anonymous):

620 000 meters = 620 kilometers

OpenStudy (anonymous):

\[\frac{1km}{km?}=\frac{1000m}{620,000m}\]cross and multiply \[1km \times620,000m=1000m \times .km\] \[\frac{620,000}{1000}=km\]\[620km=620,000m\]